Find the present values of the following cash flow streams. The appropriate interest rate is 13%....

  1. Find the present values of the following cash flow streams. The appropriate interest rate is 13%. Round your answers to the nearest cent. (Hint: It is fairly easy to work this problem dealing with the individual cash flows. However, if you have a financial calculator, read the section of the manual that describes how to enter cash flows such as the ones in this problem. This will take a little time, but the investment will pay huge dividends throughout the course. Note that, when working with the calculator's cash flow register, you must enter CF0 = 0. Note also that it is quite easy to work the problem with Excel, using procedures described in the Chapter 4 Tool Kit.)
    Year Cash Stream A Cash Stream B
    1 $100 $300
    2 400 400
    3 400 400
    4 400 400
    5 300 100

    Stream A $   
    Stream B $   
  2. What is the value of each cash flow stream at a 0% interest rate? Round your answers to the nearest cent.
    Stream A $   
    Stream B $  
Answer #1

Present value of cash stream A = 100*PVF(13℅, 1 year) + 400*PVF(13%, 2 years)+400*PVF(13%, 3 years) + 400*PVF(13%, 4 Years) + 300*PVF(13%, 5 years)

= 100*0.885+400*0.783+400*0.693+400*0.613+300*0.543

=$1,087

Present value of cash stream B= 300*PVF(13℅, 1 year) + 400*PVF(13%, 2 years)+400*PVF(13%, 3 years) + 400*PVF(13%, 4 Years) + 100*PVF(13%, 5 years)

= 300*0.885+400*0.783+400*0.693+400*0.613+100*0.543

=$1,155.4

Present value of stream of cash flows at 0% interest rate

Cash Stream A = 100+400+400+400+300

= $1,600

Cash Stream B = 300+400+400+400+100

=$1,600